OpenStudy (anonymous):

integral of radical2x+1/x+1 is there any one who can help me

OpenStudy (anonymous):

 2x+1/x+1 = assume x+1=t then 2x+1=2t-1 and dx=dt which becomes integral (2t-1/t )dt split about - sign into 2 integrals and replace the value of t afterwards